A chemist has a solution of 65% pure acid and another solution of 30% pure acid. How many gallons of each will make 350 gallons of 55% pure acid solution?
Answer by TimothyLamb(4379)

x = gallons of 65% acid
y = gallons of 30% acid
---
x + y = 350
0.65x + 0.30y = 0.55*350
---
put the system of linear equations into standard form
---
x + y = 350
0.65x + 0.30y = 192.5
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = 250
y = 100
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---